OpenStudy (toxicsugar22):

Determine the thickness, in mm, of a 7.0 cm* 3.5 cm of a piece of aluminum foil if it has a mass of 0.1260 g.

OpenStudy (anonymous):

We know that density is defined as\[\rho = {\rm mass \over \rm volume}\] We can find the thickness \(t\) as\[\rho = {m \over l \times w \times t}\] where \(l\) is the length and \(w\) is the width.

OpenStudy (toxicsugar22):

24.5 right

OpenStudy (anonymous):

That doesn't seem right. What value for density are you using?

OpenStudy (anonymous):

Try \(2.70 {\rm g \over cm^3}\)

OpenStudy (mathstudent55):

The density, d, is mass, m, divided by volume, v. \(d = \dfrac{m}{v} \) That means the volume is \(v = \dfrac{m}{d} \) Using the mass of the piece of aluminum foil and the density of aluminum, can you find the volume of this piece of aluminum foil?

OpenStudy (toxicsugar22):

is it 24.5

OpenStudy (mathstudent55):

No.

OpenStudy (toxicsugar22):

then can you show me how

OpenStudy (toxicsugar22):

as it is very hard to communicate through online. Can you just give me the solutions and anser without interuptions

OpenStudy (toxicsugar22):

and then ifi have any more questions. i will tellme teacher at school or at theend i will ask you any questions i have

OpenStudy (mathstudent55):

Use the equation I gave above: \(v = \dfrac{m}{d} \) \(v = \dfrac{0.1260~g}{2.7 \frac{g}{cm^3}} \) \(v = 0.0466666 ~cm^3\)

OpenStudy (mathstudent55):

The next step is to find the thickness. \(v = lwh\) \(h = \dfrac{v}{lw} \) \(h = \dfrac{0.046666~cm^3}{(7.0~cm)(3.5~cm)} \) This will give you the height (thickness) of the piece of aluminum foil in cm. Then you need to convert to mm.